How do you solve this algebra problem?
6=1-2n+5

Answer by checkley77(12844)

6=1-2n+5 combine the integers.
6-1-5=-2n
6-6=-2n
0=-2n divide bith sides by 2
0=-n multiply both sides by -1.
n=0 ans.
Proof:
6=1-2*0+5
6=1+5
6=6